# MINT: Multi-Vector Search Index Tuning

### v1

@@ 3.2. NP-Hardness of Query Planning
Problem 3 (Set Cover $(\mathscr{U},\mathscr{S},\mathscr{t})$).
Given a universe set $\mathscr{U}=\{1,2,3,...,\mathscr{n}\}$ and a set of subsets $\mathscr{S}=\{\mathscr{s}_{i}\}$ each of whose elements $\mathscr{s}_{i}$ is a subset of $\mathscr{U}$, a set $\mathscr{C}\subseteq\mathscr{S}$ is a set cover if the union of elements of $\mathscr{C}$ equals $\mathscr{U}$. Find a set cover $\mathscr{C}$ such that $|\mathscr{C}|\leq\mathscr{t}$.
@@
Intuition: We reduce the Set Cover problem to Problem 2 by making the universe set $\mathscr{U}$ be ground truth, and making each index $x_{i}$ “return $\mathscr{s}_{i}$ in limited cost” if and only if $ek_{i}=|\mathscr{U}|=\mathscr{n}$. More specifically, when $ek_{i}<\mathscr{n}$, design $res(q,x_{i},ek_{i})$ to make index $x_{i}$ return nothing in $\mathscr{s}_{i}$; when $ek_{i}>\mathscr{n}$, design $cost_{idx}(q,x_{i},ek_{i})$ to make the cost huge.
@@
(Sketch) We first reduce an arbitrary Set Cover problem to Problem 2 in polynomial time. Given $(\mathscr{U},\mathscr{S},\mathscr{t})$ in Set Cover, let $gt(q)=\mathscr{U}$ and $q.dim=|\mathscr{S}|=\mathscr{d}$, map each $\mathscr{s}_{i}\in\mathscr{S}$ to an index $x_{i}$ with $x_{i}.dim=1$, and define
$cost_{idx}(q,x_{i},ek_{i})=\left\{\begin{aligned} 0&\quad\text{when}~{}ek_{i}=0,\\ \mathscr{n}&\quad\text{when}~{}1\leq ek_{i}\leq\mathscr{n}-1,\\ 2\mathscr{n}+1&\quad\text{when}~{}ek_{i}=\mathscr{n},\\ \infty&\quad\text{when}~{}ek_{i}\geq\mathscr{n}+1.\end{aligned}\right.$
@@
$\small\text{a set of $ek_{i}$ dummy IDs}&\quad\text{when}~{}0\leq ek_{i}\leq\mathscr{n}-1,\\ \small\mathscr{s}_{i}\text{~{}and~{}}(ek_{i}-|\mathscr{s}_{i}|)\text{~{}dummy IDs}&\quad\text{when}~{}ek_{i}\geq\mathscr{n}.\end{aligned}\right.$
@@
Next, we map each Set Cover solution $\mathscr{C}$ to a Query Planning solution $EK$ and vice versa. Formally, we show that a Set Cover problem $(\mathscr{U},\mathscr{S},\mathscr{t})$ has a solution $\mathscr{C}$ if and only if Problem 2 has a solution $EK$ satisfying $recall(q,X,EK)=100\%$ and $cost_{plan}(q,X,EK)\leq(2\mathscr{n}+1)\cdot\mathscr{t}+\mathscr{d}\mathscr{n}\cdot\mathscr{t}$.
(A) We first prove the mapping from $\mathscr{C}$ to $EK$.
Formally, given a Set Cover solution $\mathscr{C}$ where $|\mathscr{C}|\leq\mathscr{t}$, a solution $EK$ exists for Problem 2 satisfying $recall(q,X,EK)=100\%$ and $cost_{plan}(q,X,EK)\leq(2\mathscr{n}+1)\cdot\mathscr{t}+\mathscr{d}\mathscr{n}\cdot\mathscr{t}$.
For each $\mathscr{s}_{i}\in\mathscr{S}$, let $ek_{i}=\mathscr{n}$ if $\mathscr{s}_{i}\in\mathscr{C}$, otherwise $ek_{i}=0$.
  • $recall(q,X,EK)=100\%$ because $\mathscr{C}$ is a set cover that makes $\cup_{i=1}^{|X|}res(q,x_{i},ek_{i})=\cup_{\mathscr{s}_{i}\in\mathscr{C}}(\mathscr{s}_{i}\text{~{}and dummy IDs})\supseteq\mathscr{U}$.
  • $cost_{plan}(q,X,EK)\leq(2\mathscr{n}+1)\cdot\mathscr{t}+\mathscr{d}\mathscr{n}\cdot\mathscr{t}$ because by Equation 4:
@@
    $=$ $\quad\textstyle\sum_{i=1}^{|X|}\Big{(}cost_{idx}(q,x_{i},ek_{i})+\mathscr{d}\cdot ek_{i}\Big{)}$
    For those $ek_{i}=0$, the cost is 0. For those $ek_{i}=n$, the cost is $(2\mathscr{n}+1+\mathscr{d}\mathscr{n})$. So $cost_{plan}(q,X,EK)=|\mathscr{C}|\cdot(2\mathscr{n}+1+\mathscr{d}\mathscr{n})\leq\mathscr{t}\cdot(2\mathscr{n}+1+\mathscr{d}\mathscr{n})$
(B) We now prove the mapping from $EK$ to $\mathscr{C}$.
Formally, given a solution $EK$ satisfying $recall(q,X,EK)=100\%$ and $cost_{plan}(q,X,EK)\leq(2\mathscr{n}+1)\cdot\mathscr{t}+\mathscr{d}\mathscr{n}\cdot\mathscr{t}$, a set cover $\mathscr{C}$ exists and $|\mathscr{C}|\leq\mathscr{t}$.
Let $\mathscr{C}=\{\mathscr{s}_{i}~{}|~{}ek_{i}\geq\mathscr{n}\}$.
  • We prove $\mathscr{C}$ is a set cover: $recall(q,X,EK)=100\%$ means $\cup_{\mathscr{s}_{i}\in\mathscr{C}}=\mathscr{U}$ because dummy IDs do not impact recall.
  • Prove $|\mathscr{C}|\leq\mathscr{t}$:
    By Equation 4, $cost_{plan}(q,X,EK)=\sum_{i=1}^{|X|}cost_{idx}(q,x_{i},ek_{i})\allowbreak+\allowbreak cost_{rerank}(q,EK)=\sum_{i=1}^{|X|}\Big{(}cost_{idx}(q,x_{i},ek_{i})+\mathscr{d}\cdot ek_{i}\Big{)}$.
    Since $cost_{plan}(q,X,EK)\leq(2\mathscr{n}+1)\cdot\mathscr{t}+\mathscr{d}\mathscr{n}\cdot\mathscr{t}$, there must be no $ek_{i}\geq\mathscr{n}+1$ that makes $cost_{idx}(q,x_{i},ek_{i})=\infty$.
@@
    Because $\sum_{1\leq ek_{i}\leq\mathscr{n}-1}(\mathscr{n}+\mathscr{d}\cdot ek_{i})\geq 0$,
    $\therefore|\mathscr{C}|(2\mathscr{n}+1+\mathscr{d}\cdot\mathscr{n})\leq cost_{plan}(q,X,EK)\leq(2\mathscr{n}+1)\cdot\mathscr{t}+\mathscr{d}\mathscr{n}\cdot\mathscr{t}$
    $\therefore|\mathscr{C}|\leq\mathscr{t}$.
We can provide an example in which the designed $cost_{idx}(q,x_{i},ek_{i})$ and $res(q,x_{i},ek_{i})$ are real. The example is an HNSW index with one layer graph. Given $ek_{i}$, the HNSW search algorithm (Malkov and Yashunin, 2018) starts from a fixed entry point and explores neighbors to update the top-$ek_{i}$ items. The search stops when the new neighbors cannot make top-$ek_{i}$ items better (i.e., top-$ek_{i}$ items converge). That means, even when we increase $ek_{i}$ by $1$, more than one neighbor can be explored. As we will see, when $ek_{i}\leq\mathscr{n}-1$, the search stops after scanning $\mathscr{n}$ items. When $ek_{i}=\mathscr{n}$, the search stops after scanning $(2\mathscr{n}+1)$ items. When $ek_{i}\geq\mathscr{n}+1$, the search has to scan a large number of items (i.e., all items in the database).
Figure 3. Example HNSW index with one layer for $\mathscr{n}=3$. IDs $\{1,2,3\}=\mathscr{U}$ in Problem 3, while $\text{IDs}\geq 91$ are dummy IDs. When $21\leq ek_{i}\leq\mathscr{n}-1=$ 2, $\mathscr{n}=3$ items in the first row are scanned. When $ek_{i}=\mathscr{n}=3$, $(2\mathscr{n}+1)=7$ items in the first and seconds rows are scanned. When $ek_{i}\geq\mathscr{n}+1$ = 4, all items in the database are scanned.
@@
Figure 3 is an HNSW index with one layer graph. Let $\mathscr{n}=3$. IDs $1$ to $3$ are in the universe set $\mathscr{U}$ in Problem 3, while $\text{IDs}\geq 91$ are dummy IDs.
  • When $ek_{i}=0$, the index is not scanned. $cost_{idx}(q,x_{i},ek_{i})=0$ and $res(q,x_{i},ek_{i})=\varnothing$
  • When $21\leq ek_{i}\leq\mathscr{n}-1=$ 2, the search explores items $91$ and $92$ and stops because $92$’s new neighbor $93$ does not improve the top items. $cost_{idx}(q,x_{i},ek_{i})=\mathscr{n}=3$ because the scores of $\{91,92,93\}$ are computed. $res(q,x_{i},ek_{i})$ contains only dummy IDs.
  • When $ek_{i}=\mathscr{n}=3$, the search further explores $93$’s neighbor to find better items $\{1,2,3\}$, and stops because item $1$’s neighbor $94$ does not improve top-3. So $cost_{idx}(q,x_{i},ek_{i})=2\mathscr{n}+1=7$ for $\{1,2,3,91,92,93,94\}$. $res(q,x_{i},ek_{i})=\{1,2,3\}$.
  • When $ek_{i}\geq\mathscr{n}+1=4$, the search further explores $94$’s neighbor to find $95$, and then has to explore all remaining items like $96$, $97$, etc. So $cost_{idx}(q,x_{i},ek_{i})\approx\infty$. $res(q,x_{i},ek_{i})=\{1,2,3\}\cup\{\text{dummy IDs}\}$.
3.3. Dynamic Programming Solution for Query Planning
3.3.1. Search and Dynamic Programming Algorithms
Although Problem 2 is NP-hard, we can address it by leveraging the limited $k$ and $EK$ in practice. In real-world applications, $k$ is usually around $10$ to a few hundred (Aumüller et al., 2020; Simhadri et al., 2022, 2024). Assume for a moment that we can get the ground truth $gt(q)$ and $cost_{idx}(q,x_{i},ek_{i})$; we will relax this assumption and discuss how to estimate them in Section 3.3.2. For each item in the ground truth $gt(q)$, we can identify the item’s ranking in each index $x_{i}$ as a relevant $ek_{i}$. In other words, for $x_{i}$, the number of relevant $ek_{i}$ is limited to $k$.
Figure 4. Example of relevant $ek$. The number of each relevant $ek$ is limited to $k=3$. The search space of $ek^{1}$ and $ek^{3}$ is $\{0,6,31,183\}\times\{0,8,116,230\}$, which has $(k+1)(k+1)=16$ combinations. A possible plan is $ek^{1}=31$ and $ek^{3}=8$, which covers $r_{1}$, $r_{2}$, and $r_{3}$.
@@ 3.3.1. Search and Dynamic Programming Algorithms
Take $ek^{1}$ as an example. For the $j$-th item in the ground truth $gt(q^{1,3})$ ranked by $\sum_{i\in\{1,3\}}score(q^{1,3}.\textbf{v}_{i},r.\textbf{v}_{i})$, we can find its ranking $ek^{1}\_j$ in $x^{1}$ by $score(q^{1,3}.\textbf{v}_{1},r.\textbf{v}_{1})$. For instance, the 1st item in $gt(q^{1,3})$ may have $ek^{1}\_1=31$ in $x^{1}$, the 2nd item in $gt(q^{1,3})$ may have $ek^{1}\_2=6$, and so on. If no item in $gt(q^{1,3})$ ranks between $7$ and $30$, we do not care $7\leq ek^{1}\leq 30$ because, compared to $ek^{1}=6$, $ek^{1}$ between $7$ and $30$ do not cover more items in $gt(q^{1,3})$. In other words, only the $3$ $ek^{1}\_j$ are relevant.
Similarly, we can find the $3$ rankings $ek^{3}\_j$ in $x^{3}$ by $score(q^{1,3}.\textbf{v}_{3},r.\textbf{v}_{3})$.
Then, we add $ek^{1}\_0=0$ and $ek^{3}\_0=0$ indicating the corresponding index may be skipped.
Finally, we only focus on search space $\{ek^{1}\_j\}\times\{ek^{3}\_j\}$ whose size is $(k+1)\cdot(k+1)$.
Because of the limited number of relevant $ek_{i}$, we can enumerate the combinations of $ek_{i}$ to find the best plan. Given indexes $X$, the search space of $EK$ is $(k+1)^{|X|}$. The search algorithm’s space complexity is $O(k\cdot|X|)$ for pre-computing the relevant $ek_{i}$. The time complexity is $O(k\cdot\sum_{x_{i}\in X}x_{i}.dim+|X|k\log k+k^{|X|})$ where $k\cdot\sum_{x_{i}\in X}x_{i}.dim$ is to compute the scores on $x_{i}$, $|X|k\log k$ is to sort $ek_{i}\_j$ for each $x_{i}$, and $k^{|X|}$ is to enumerate the combinations of $ek_{i}$ and incrementally counting the covered items.
Input: Query $q$; Top $k$; Candidate indexes $X$; Recall threshold $\theta_{recall}$; Ground truth $gt(q)$; Cost function $cost_{plan}(q,X,EK)$.
Output: Query Plan $(X,EK)$
1 foreach $1\leq i\leq|X|$ do
2       $docPairs\leftarrow$$\{(ID:j,Rank:ek_{i}\_j)|1\leq j\leq k\}\cup\{(ID:null,Rank:0)\}$
3       $ek_{i}\_relevant\leftarrow$ sort $docPairs$ by Rank
4
5$(bestEK,bestCost)\leftarrow(null,\infty)$
Search ($1,\varnothing$)
  // Defined below
6 return $(X,bestEK)$
7
8
9Function Search ($𝑜𝑣𝑒𝑟i,c$)
       /* Assuming $|X|\geq 2$/
10       if $i<|X|-1$ then
11             foreach $0\leq ek_{i}\leq k$ do
[…]
13                   Search ($𝑜𝑣𝑒𝑟i+1,c$)
14
15
16      else
             /* $i=|X|-1$/
17             $ek_{|X|}\leftarrow k$
18             $cvr_{|X|}\leftarrow\{1\leq j\leq k\}$
19             foreach $0\leq ek_{i}\leq k$ do
[…]
21                   Get the minimum $ek_{|X|}$ and its corresponding $cvr_{|X|}$ satisfying $|cover\cup cvr_{|X|}|\geq\theta_{recall}\cdot k$
22                   Update $(bestEK,bestCost)$ if the cost of the current $\{ek_{i^{\prime}}|1\leq i^{\prime}\leq|X|\}$ is lower
23
24
25
Algorithm 1 Search
When $|X|\geq 2$, we can improve this search algorithm by computing the last $ek_{|X|}$ from the other $ek_{i}$ where $11\leq i\leq|X|-$ 1. For example, when $|X|=2$, as we increase $ek_{1}$, the smallest $ek_{2}$ to meet the recall threshold must be decreasing. So we can initialize $ek_{2}=k$ and keep decreasing $ek_{2}$ when we increase $ek_{1}$. This optimization removes the loop of $ek_{|X|}$ and brings the time complexity down to $O(k\cdot\sum_{x_{i}\in X}x_{i}.dim+|X|k\log k+k^{|X|\textbf{-1}})$.
Algorithm 1 is the pseudo code of the search algorithm. Line 1 to Line 1 compute $ek_{i}\_j$ for the $j$-th item in the ground truth $gt(q)$, and $ek_{i}\_0=0$ means index $x_{i}$ can be skipped. In the $Search(...)$ function, when $i<|X|-1$, we enumerate $ek_{i}$ and update the covered items (Line 1 to Line 1). When $i=|X|-1$, we increase $ek_{i}$ (which is $ek_{|X|-1}$) and decrease $ek_{|X|}$ to the minimum to cover $\theta_{recall}\cdot k$ items (Line 1 to Line 1).
@@ 13                   Search ($𝑜𝑣𝑒𝑟i+1,c$)
1 foreach $1\leq i\leq|X|$ do
2       $docPairs\leftarrow$$\{(ID:j,Rank:ek_{i}\_j)|1\leq j\leq k\}\cup\{(ID:null,Rank:0)\}$
3       $ek_{i}\_relevant\leftarrow$ sort $docPairs$ by Rank
4
5
/* $DP(i,cover)$ for cost; $EK(i,cover)$ for $EK$ values. */
6 foreach $cover\in\mathcal{P}(gt_{sp}(q))$ do
7       $DP(1,cover)\leftarrow cost_{cover}(cover,x_{1})$
       $EK(1,cover)\leftarrow\{ek_{1}:\max_{j\in cover}ek_{1}\_j\}$
@@ 7       $DP(1,cover)\leftarrow cost_{cover}(cover,x_{1})$
8
9
10foreach $2\leq i\leq|X|$ do
11       foreach $cover\in\mathcal{P}(gt_{sp}(q))$ do
12             $DP(i,cover)\leftarrow\infty$
13             foreach $𝑜𝑣𝑒𝑟cvr\subseteq c$ do
14                   if $DP(i,cover)>DP(i-1,cover-cvr)+cost_{cover}(cvr,x_{i})$ then
15                         $DP(i,cover)\leftarrow DP(i-1,cover-cvr)+cost_{cover}(cvr,x_{i})$
16                         $EK(i,cover)\leftarrow EK(i-1,cover-cvr)\cup\{ek_{i}:\max_{j\in cvr}ek_{i}\_j\}$
17
18
19
20
[…]
22 $bestEK\leftarrow EK(|X|,bestCover)$
23 return $(X,bestEK)$
24
@@ 16                         $EK(i,cover)\leftarrow EK(i-1,cover-cvr)\cup\{ek_{i}:\max_{j\in cvr}ek_{i}\_j\}$
Algorithm 1 may not scale well due to the $k^{|X|-1}$ in its time complexity. So we further optimize it by proposing a Dynamic Programming algorithm without $|X|$ as an exponent.

### v2

@@ III-B NP-Hardness of Query Planning
Problem 3 (Set Cover $(\mathtt{U},\mathtt{S},\mathtt{t})$).
Given a universe set $\mathtt{U}=\{1,2,3,...,\mathtt{n}\}$ and a set of subsets $\mathtt{S}=\{\mathtt{s}_{i}\}$ each of whose elements $\mathtt{s}_{i}$ is a subset of $\mathtt{U}$, a set $\mathtt{C}\subseteq\mathtt{S}$ is a set cover if the union of elements of $\mathtt{C}$ equals $\mathtt{U}$. Find a set cover $\mathtt{C}$ such that $|\mathtt{C}|\leq\mathtt{t}$.
@@
Intuition: We reduce the Set Cover problem to Problem 2 by making the universe set $\mathtt{U}$ be ground truth, and making each index $x_{i}$ “return $\mathtt{s}_{i}$ in limited cost” if and only if $ek_{i}=|\mathtt{U}|=\mathtt{n}$. More specifically, when $ek_{i}<\mathtt{n}$, design $res(q,x_{i},ek_{i})$ to make index $x_{i}$ return nothing in $\mathtt{s}_{i}$; when $ek_{i}>\mathtt{n}$, design $cost_{idx}(q,x_{i},ek_{i})$ to make the cost huge.
@@
(Sketch) We first reduce an arbitrary Set Cover problem to Problem 2 in polynomial time. Given $(\mathtt{U},\mathtt{S},\mathtt{t})$ in Set Cover, let $gt(q)=\mathtt{U}$ and $q.dim=|\mathtt{S}|=\mathtt{d}$, map each $\mathtt{s}_{i}\in\mathtt{S}$ to an index $x_{i}$ with $x_{i}.dim=1$, and define
$cost_{idx}(q,x_{i},ek_{i})=\left\{\begin{aligned} 0&\quad\text{when}~ek_{i}=0,\\ \mathtt{n}&\quad\text{when}~1\leq ek_{i}\leq\mathtt{n}-1,\\ 2\mathtt{n}+1&\quad\text{when}~ek_{i}=\mathtt{n},\\ \infty&\quad\text{when}~ek_{i}\geq\mathtt{n}+1.\end{aligned}\right$.
@@
[…]
@@
Next, we map each Set Cover solution $\mathtt{C}$ to a Query Planning solution $EK$ and vice versa. Formally, we show that a Set Cover problem $(\mathtt{U},\mathtt{S},\mathtt{t})$ has a solution $\mathtt{C}$ if and only if Problem 2 has a solution $EK$ satisfying $recall(q,X,EK)=100\%$ and $cost_{plan}(q,X,EK)\leq(2\mathtt{n}+1)\cdot\mathtt{t}+\mathtt{d}\mathtt{n}\cdot\mathtt{t}$.
(A) We first prove the mapping from $\mathtt{C}$ to $EK$.
Formally, given a Set Cover solution $\mathtt{C}$ where $|\mathtt{C}|\leq\mathtt{t}$, a solution $EK$ exists for Problem 2 satisfying $recall(q,X,EK)=100\%$ and $cost_{plan}(q,X,EK)\leq(2\mathtt{n}+1)\cdot\mathtt{t}+\mathtt{d}\mathtt{n}\cdot\mathtt{t}$.
For each $\mathtt{s}_{i}\in\mathtt{S}$, let $ek_{i}=\mathtt{n}$ if $\mathtt{s}_{i}\in\mathtt{C}$, otherwise $ek_{i}=0$.
  • $recall(q,X,EK)=100\%$ because $\mathtt{C}$ is a set cover that makes $\cup_{i=1}^{|X|}res(q,x_{i},ek_{i})=\cup_{\mathtt{s}_{i}\in\mathtt{C}}(\mathtt{s}_{i}\text{~and dummy IDs})\supseteq\mathtt{U}$.
  • $cost_{plan}(q,X,EK)\leq(2\mathtt{n}+1)\cdot\mathtt{t}+\mathtt{d}\mathtt{n}\cdot\mathtt{t}$ because by Equation 4:
@@
    $=$ $\quad\textstyle\sum_{i=1}^{|X|}\Big(cost_{idx}(q,x_{i},ek_{i})+\mathtt{d}\cdot ek_{i}\Big)$
    For those $ek_{i}=0$, the cost is 0. For those $ek_{i}=n$, the cost is $(2\mathtt{n}+1+\mathtt{d}\mathtt{n})$. So $cost_{plan}(q,X,EK)=|\mathtt{C}|\cdot(2\mathtt{n}+1+\mathtt{d}\mathtt{n})\leq\mathtt{t}\cdot(2\mathtt{n}+1+\mathtt{d}\mathtt{n})$
(B) We now prove the mapping from $EK$ to $\mathtt{C}$.
Formally, given a solution $EK$ satisfying $recall(q,X,EK)=100\%$ and $cost_{plan}(q,X,EK)\leq(2\mathtt{n}+1)\cdot\mathtt{t}+\mathtt{d}\mathtt{n}\cdot\mathtt{t}$, a set cover $\mathtt{C}$ exists and $|\mathtt{C}|\leq\mathtt{t}$.
Let $\mathtt{C}=\{\mathtt{s}_{i}~|~ek_{i}\geq\mathtt{n}\}$.
  • We prove $\mathtt{C}$ is a set cover: $recall(q,X,EK)=100\%$ means $\cup_{\mathtt{s}_{i}\in\mathtt{C}}=\mathtt{U}$ because dummy IDs do not impact recall.
  • Prove $|\mathtt{C}|\leq\mathtt{t}$:
    By Equation 4, $cost_{plan}(q,X,EK)=\sum_{i=1}^{|X|}cost_{idx}(q,x_{i},ek_{i})\allowbreak+\allowbreak cost_{rerank}(q,EK)=\sum_{i=1}^{|X|}\Big(cost_{idx}(q,x_{i},ek_{i})+\mathtt{d}\cdot ek_{i}\Big)$.
    Since $cost_{plan}(q,X,EK)\leq(2\mathtt{n}+1)\cdot\mathtt{t}+\mathtt{d}\mathtt{n}\cdot\mathtt{t}$, there must be no $ek_{i}\geq\mathtt{n}+1$ that makes $cost_{idx}(q,x_{i},ek_{i})=\infty$.
@@
    Because $\sum_{1\leq ek_{i}\leq\mathtt{n}-1}(\mathtt{n}+\mathtt{d}\cdot ek_{i})\geq 0$,
    $\therefore|\mathtt{C}|(2\mathtt{n}+1+\mathtt{d}\cdot\mathtt{n})\leq cost_{plan}(q,X,EK)\leq(2\mathtt{n}+1)\cdot\mathtt{t}+\mathtt{d}\mathtt{n}\cdot\mathtt{t}$
    $\therefore|\mathtt{C}|\leq\mathtt{t}$.
We provide an example in which the designed $cost_{idx}(q,x_{i},ek_{i})$ and $res(q,x_{i},ek_{i})$ are real. The example is an HNSW index with one layer graph. Given $ek_{i}$, the HNSW search algorithm [55] starts from a fixed entry point and explores neighbors to update the top-$ek_{i}$ items. The search stops when the new neighbors cannot make top-$ek_{i}$ items better (i.e., top-$ek_{i}$ items converge). That means, even when we increase $ek_{i}$ by $1$, more than one neighbor can be explored. As we will see, when $ek_{i}\leq\mathtt{n}-1$, the search stops after scanning $\mathtt{n}$ items. When $ek_{i}=\mathtt{n}$, the search stops after scanning $(2\mathtt{n}+1)$ items. When $ek_{i}\geq\mathtt{n}+1$, the search has to scan a large number of items (i.e., all items in the database).
Figure 3: Example HNSW index with one layer for $\mathtt{n}=3$. IDs $\{1,2,3\}=\mathtt{U}$ in Problem 3, while $\text{IDs}\geq 91$ are dummy IDs. When $1\leq ek_{i}\leq\mathtt{n}-1=2$, $\mathtt{n}=3$ items in the first row are scanned. When $ek_{i}=\mathtt{n}=3$, $(2\mathtt{n}+1)=7$ items in the first and seconds rows are scanned. When $ek_{i}\geq\mathtt{n}+1$ = 4, all items in the database are scanned.
@@
Figure 3 is an HNSW index with one layer graph. Let $\mathtt{n}=3$. IDs $1$ to $3$ are in the universe set $\mathtt{U}$ in Problem 3, while $\text{IDs}\geq 91$ are dummy IDs.
  • When $ek_{i}=0$, the index is not scanned. $cost_{idx}(q,x_{i},ek_{i})=0$ and $res(q,x_{i},ek_{i})=\varnothing$.
  • When $1\leq ek_{i}\leq\mathtt{n}-1=2$, the search explores items $91$ and $92$ and stops because $92$’s new neighbor $93$ does not improve the top items. $cost_{idx}(q,x_{i},ek_{i})=\mathtt{n}=3$ because the scores of $\{91,92,93\}$ are computed. $res(q,x_{i},ek_{i})$ contains only dummy IDs.
  • When $ek_{i}=\mathtt{n}=3$, the search further explores $93$’s neighbor to find better items $\{1,2,3\}$, and stops because item $1$’s neighbor $94$ does not improve top-3. So $cost_{idx}(q,x_{i},ek_{i})=2\mathtt{n}+1=7$ for $\{1,2,3,91,92,93,94\}$. $res(q,x_{i},ek_{i})=\{1,2,3\}$.
  • When $ek_{i}\geq\mathtt{n}+1=4$, the search further explores $94$’s neighbor to find $95$, and then has to explore all remaining items like $96$, $97$, etc. So $cost_{idx}(q,x_{i},ek_{i})\approx\infty$. $res(q,x_{i},ek_{i})=\{1,2,3\}\cup\{\text{dummy IDs}\}$.
III-C Dynamic Programming Solution for Query Planning
III-C1 Search and Dynamic Programming Algorithms
Although Problem 2 is NP-hard, we can address it by leveraging the limited $k$ and $EK$ in practice. In real-world applications, $k$ is usually around $10$ to a few hundred [5, 73, 72]. Assume for a moment that we can get the ground truth $gt(q)$ and $cost_{idx}(q,x_{i},ek_{i})$; we will relax this assumption and discuss how to estimate them in Section III-C2. For each item in the ground truth $gt(q)$, we can identify the item’s ranking in each index $x_{i}$ as a relevant $ek_{i}$. In other words, for $x_{i}$, the number of relevant $ek_{i}$ is limited to $k$.
Figure 4: Example of relevant $ek$. The number of each relevant $ek$ is limited to $k=3$. The search space of $ek^{1}$ and $ek^{3}$ is $\{0,6,31,183\}\times\{0,8,116,230\}$, which has $(k+1)(k+1)=16$ combinations. A possible plan is $ek^{1}=31$ and $ek^{3}=8$, which covers $r_{1}$, $r_{2}$, and $r_{3}$.
@@ III-C1 Search and Dynamic Programming Algorithms
Take $ek^{1}$ as an example. For the $j$-th item in the ground truth $gt(q^{1,3})$ ranked by $\sum_{i\in\{1,3\}}score(q^{1,3}.\textbf{v}_{i},r.\textbf{v}_{i})$, we can find its ranking $ek^{1}.j$ in $x^{1}$ by $score(q^{1,3}.\textbf{v}_{1},r.\textbf{v}_{1})$. For instance, the 1st item in $gt(q^{1,3})$ may have $ek^{1}.1=31$ in $x^{1}$, the 2nd item in $gt(q^{1,3})$ may have $ek^{1}.2=6$, and so on. If no item in $gt(q^{1,3})$ ranks between $7$ and $30$, we do not care $7\leq ek^{1}\leq 30$ because, compared to $ek^{1}=6$, $ek^{1}$ between $7$ and $30$ do not cover more items in $gt(q^{1,3})$. In other words, only the 33 $ek^{1}.j$ are relevant.
Similarly, we can find the $3$ rankings $ek^{3}.j$ in $x^{3}$ by $score(q^{1,3}.\textbf{v}_{3},r.\textbf{v}_{3})$.
Then, we add $ek^{1}.0=0$ and $ek^{3}.0=0$ indicating the corresponding index may be skipped.
Finally, we only focus on search space $\{ek^{1}.j\}\times\{ek^{3}.j\}$ whose size is $(k+1)\cdot(k+1)$.
Because of the limited number of relevant $ek_{i}$, we can enumerate the combinations of $ek_{i}$ to find the best plan (detail in [88]). Given indexes $X$, the search space of $EK$ is $(k+1)^{|X|}$. The search algorithm’s space complexity is $O(k\cdot|X|)$ for pre-computing the relevant $ek_{i}$. The time complexity is $O(k\cdot\sum_{x_{i}\in X}x_{i}.dim+|X|k\log k+k^{|X|})$ where $k\cdot\sum_{x_{i}\in X}x_{i}.dim$ is to compute the scores on $x_{i}$, $|X|k\log k$ is to sort $ek_{i}.j$ for each $x_{i}$, and $k^{|X|}$ is to enumerate the combinations of $ek_{i}$ and incrementally counting the covered items. The $k^{|X|}$ term limits its scalability. So we further optimize it by proposing a Dynamic Programming algorithm without $|X|$ as an exponent.
@@
foreach $1\leq i\leq|X|$ do
  $docPairs\leftarrow$ $\{(ID:j,Rank:ek_{i}.j)|1\leq j\leq k\}\cup\{(ID:null,Rank:0)\}$
  $ek_{i}.relevant\leftarrow$ sort docPairsdocPairs by Rank
end foreach
/* $DP(i,cover)$ for cost; $EK(i,cover)$ for EKEK values. */
foreach $cover\in\mathcal{P}(gt_{sp}(q))$ do
  $DP(1,cover)\leftarrow cost_{cover}(cover,x_{1})$
  $EK(1,cover)\leftarrow\{ek_{1}:\max_{j\in cover}ek_{1}.j\}$
@@
end foreach
foreach $2\leq i\leq|X|$ do
  foreach $cover\in\mathcal{P}(gt_{sp}(q))$ do
     $DP(i,cover)\leftarrow\infty$
     foreach $cvr\subseteq cover$ do
        if $DP(i,cover)>DP(i-1,cover-cvr)+cost_{cover}(cvr,x_{i})$ then
           $DP(i,cover)\leftarrow DP(i-1,cover-cvr)+cost_{cover}(cvr,x_{i})$
           $EK(i,cover)\leftarrow EK(i-1,cover-cvr)\cup\{ek_{i}:\max_{j\in cvr}ek_{i}.j\}$
        end if
     end foreach
  end foreach
end foreach
$bestCover\leftarrow arg\,min_{|cover|\geq k^{\prime}\cdot\theta_{recall}}DP(|X|,cover)$
$bestEK\leftarrow EK(|X|,bestCover)$
return $(X,bestEK)$